OpenStudy (anonymous):

The radius of a circle measures twenty centimeters. What is the length of its diameter? 5 cm 10 cm 20 cm 40 cm

OpenStudy (acxbox22):

radius=diameter/2 r=d/2 20=d/2 d=40cm

OpenStudy (anonymous):

ok. Thanks!

jimthompson5910 (jim_thompson5910):

I would write it as diameter = 2*(radius) d = 2*r d = 2*20 d = 40 but either way, you get the diameter to be 40 cm